# StochasticBarrier.jl: A Toolbox for Stochastic Barrier Function Synthesis

### v1

@@ 2. Theoretical Framework
We provide an overview of the SBF theoretical framework.
@@ 2.2.2. PWC Method
Work (Mazouz et al., 2024a) introduces three methods for synthesizing PWC SBFs: (i) an exact duality-based approach formulated as a zero-duality-gap Linear Program (LP), (ii) an exact Counter-Example Guided Inductive Synthesis (CEGIS) method that alternates between generating candidate barriers and counter-examples via two LPs, and (iii) a gradient descent (GD) method offering improved scalability and memory efficiency. StochasticBarrier.jl implements all three methods for System (1) with linear and PWA inclusion dynamics.
@@ 4. Benchmark Studies
Figure 1. Polynomial systems benchmarking of StochasticBarrier.jl vs StochasticBarrierFunctions (Santoyo et al., 2021) and PRoTECT (Wooding et al., 2024). Plots depict optimization time $\tau(s)$ and probability of safety $P_{s}$. If data points are not plotted beyond a certain barrier degree, it means the algorithm ran out of memory.
@@ Appendix A Benchmarking Setup
StochasticBarrier.jl is also equipped with an independent interface to easily run benchmarks. Fig. 3 depicts the general structure of the interface. The input is a .yaml file, in which the dynamics, the sets (safe, unsafe and initial), and the SBF synthesis method are specified. The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time.
@@ B.1. Linear Systems
where $\mathbf{x}$ is the fluid heights in the two tanks and $\mathbf{w}\sim\mathcal{N}(0,10^{-4}I)$ the additive noise. The system parameters are $\tau=0.1$, $\frac{\alpha_{1}}{A_{1}}=1$, $\frac{q_{1}}{A_{1}}=4.5$, $\frac{\alpha_{1}}{A_{2}}=1$, $\frac{\alpha_{2}}{A_{2}}=1$, and $\frac{q_{o}}{A_{2}}=-3$. The domain and initial sets are $X=[1,9]^{2}$ and $X_{0}=[2.75,3.25]^{2}$, respectively.
@@ B.2. Polynomial Systems
where $\mathbf{x}_{k}$ is the room temperature, the ambient temperature $T_{e}=-15^{\circ}$C, and the heater temperature $T_{h}=45^{\circ}$C. Conduction factors are given as $\theta=0.145,\beta=0.6$, and $\nu=-0.0120155\mathbf{x}_{k}+0.8$. The system evolves over the state space $X=[1,50]$ with the initial set $X_{0}=[19.5,20]$, and unsafe regions are $X_{u}=[1,17]\cup[23,50]$. The noise parameter is $\mathbf{w}\sim\mathcal{N}(0,\cdot 10^{-2}I)$.
@@ Appendix D Table Results
Linear 2D Contraction Map 1 2 23.56 1.00 0.00 0.00 0.58 $1.7e^{-2}$ $1.8e^{-2}$ 0.80 0.01 $1.7e^{-2}$ $1.8e^{-2}$ 0.80
4 74.86 1.00 0.00 0.00 1.51 $3.7e^{-3}$ $5.7e^{-3}$ 0.94 0.02 $3.7e^{-3}$ $5.7e^{-3}$ 0.94
8 543.29 1.00 0.00 0.00 12.85 $1.1e^{-3}$ $1.8e^{-3}$ 0.98 0.05 $1.1e^{-3}$ $1.8e^{-3}$ 0.98
12 2582.17 1.00 0.00 0.00 76.88 $1.0e^{-3}$ $1.3e^{-3}$ 0.99 0.21 $1.1e^{-3}$ $1.3e^{-3}$ 0.99
24 OM $\times$ $\times$ $\times$ 2442.43 $8.3e^{-6}$ $9.5e^{-8}$ 0.99 18.08 $9.9e^{-4}$ $1.2e^{-3}$ 0.99
30 OM $\times$ $\times$ $\times$ 9569.64 $4.6e^{-7}$ $7.8e^{-9}$ 0.99 107.14 $9.6e^{-4}$ $1.2e^{-3}$ 0.99
2D Two Tank 4 90.87 1.00 0.00 0.00 2.37 $1.00$ $3.6e^{-6}$ 0.00 0.14 $8.7e^{-2}$ $4.1e^{-2}$ 0.50
6 283.36 1.00 0.00 0.00 9.48 FAILED FAILED 0.00 0.10 $2.9e^{-2}$ $7.9e^{-5}$ 0.97
8 884.79 1.00 0.00 0.00 53.60 FAILED FAILED 0.00 0.16 $6.8e^{-4}$ $4.2e^{-5}$ 1.00
3D Room Temp. 2 109.74 FAILED FAILED 0.00 1.38 FAILED FAILED 0.00 0.12 FAILED FAILED 0.00
4 OM $\times$ $\times$ $\times$ 39.02 FAILED FAILED 0.00 0.27 $9.7e^{-2}$ $7.6e^{-3}$ 0.83
6 OM $\times$ $\times$ $\times$ 2422.07 FAILED FAILED 0.00 11.08 $9.0e^{-7}$ $8.6e^{-8}$ 1.00
8 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 20.02 $8.7e^{-6}$ $2.1e^{-6}$ 1.00
12 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3716.59 $4.1e^{-8}$ $9.7e^{-9}$ 1.00
6D Quadrotor 2 OM $\times$ $\times$ $\times$ 9587.45 FAILED FAILED 0.00 0.14 $1.0e^{-1}$ $5.9e^{-4}$ 0.89
4 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 3.35 $9.0e^{-4}$ $1.0e^{-5}$ 1.00
6 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 148.65 $8.8e^{-6}$ $1.7e^{-7}$ 1.00
Polynomial 1D Thermostat 4 0.59 FAILED FAILED 0.00 0.05 0.87 $3.3e^{-3}$ 0.09
6 0.79 FAILED FAILED 0.00 0.08 $5e^{-11} 3.3$$3e^{-12}$ 1.00
8 1.13 FAILED FAILED 0.00 0.16 $4.2e^{-5}$ $8.5e^{-6}$ 1.00
2D Oscillator 6 21.21 FAILED FAILED 0.00 0.96 1.00 $1.8e^{-8}$ 0.00
8 168.79 FAILED FAILED 0.00 3.68 1.00 $2.9e^{-8}$ 0.00
12 5502.47 FAILED FAILED 0.00 77.03 $1.2e^{-7}$ $1.2e^{-8}$ 1.00

### v2

@@ 1. Introduction
1.1. Related Work
Barrier certificates are synthesized to proof safety for dynamical systems for both deterministic (Schneeberger et al., 2023; Wang et al., 2018), and stochastic systems (Prajna et al., 2002; Jagtap et al., 2020). The latter, referred to as stochastic barrier functions (SBFs), are utilized to obtain a formal lower bound on the probability of remaining safe, starting from an initial set, under a well-defined probability measure. SBFs have been used to verify safety for linear and polynomial (Santoyo et al., 2021), non-linear (Mazouz et al., 2022), and data-driven systems (Mazouz et al., 2024b).
Sum-of-Squares (SOS) optimization is a popular method for SBF synthesis (Prajna et al., 2002, 2007; Santoyo et al., 2021; Jagtap et al., 2020; Mazouz et al., 2022). Some works utilize SOSTOOLS (Prajna et al., 2002) to obtain these SBFs. The main downside of SOSTOOLS-based SBFs is that the algorithms are embedded in MATLAB, which scales poorly when bridging to SDP. PRoTect (Wooding et al., 2024), a Python based implementation, scales better, but is limited to low-degree SOS barriers and systems of polynomial class. Our toolbox, which is implemented in Julia, provides significant speedups for intensive numerical optimization tasks and memory management (Perkel and others, 2019; Coleman et al., 2021).
Related tools for the verification of stochastic systems in Julia include IntervalMDP.jl (Mathiesen et al., 2024a), which performs CUDA-accelerated value iteration for Interval Markov Decision Processes (IMDPs). JuliaReach (Bogomolov et al., 2019) performs set-based reachability for deterministic systems. A tool for Satisfiability modulo theories, Satisfiability.jl (Soroka et al., 2023), provides high-level representations for SMT formulae. Finally, ModelVerification.jl (Wei et al., 2024) verifies DNNs and safety specifications. None of these verify safety for stochastic systems using SBFs. A platform for this purpose is provided in StochasticBarrier.jl.
The remainder of the paper is organized as follows. In Section 3, we provide a detailed overview of the toolbox, with instructions and examples on setting up system models and invoking SBF synthesis engines. In Section 4, we demonstrate the performance of StochasticBarrier.jl through a series of case studies and benchmark comparisons. Finally, we conclude the paper with a summary and closing remarks in Section 5.
@@ 2. Theoretical Framework
We provide an overview of the SBF theoretical framework. An outline of the problem formulation is provided, followed by a review of the implemented methods. For more detail, we refer the reader to (Mazouz et al., 2022, 2024a).
@@ 2.2.2. PWC Method
Work (Mazouz et al., 2024a) introduces three methods for synthesizing PWC SBFs: (i) an exact duality-based approach formulated as a zero-duality-gap Linear Program (LP), (ii) an exact Counter-Example Guided Inductive Synthesis (CEGIS) method that alternates between generating candidate barriers and counter-examples via two LPs, and (iii) a gradient descent (GD) method offering improved scalability and memory efficiency. StochasticBarrier.jl implements all three methods for System (1) with linear and PWA inclusion dynamics. The toolbox internally computes the kernel bounds in Eq. (7).
@@ 4. Benchmark Studies
Figure 1. Polynomial systems benchmarking of StochasticBarrier.jl vs StochasticBarrierFunctions (Santoyo et al., 2021) and PRoTECT (Wooding et al., 2024). Plots depict optimization time $\tau(s)$ and probability of safety $P_{s}$. If data points are not plotted, it means the algorithm ran out of memory.
t Refer to caption (a) Contraction Map Refer to caption (b) Pendulum Refer to caption (c) Unicycle
@@ Appendix A Benchmarking Setup
StochasticBarrier.jl is also equipped with an independent interface to easily run benchmarks. Fig. 3 depicts the general structure of the interface. The input is a .yaml file, in which the dynamics, the sets (safe, unsafe and initial), and the SBF synthesis method are specified. The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time. The output includes the SBF $B$, values for $\eta$ and $\beta$, the lower bound on safety probability $P_{s}$, and the optimization time.
@@ B.1. Linear Systems
where $\mathbf{x}$ is the fluid heights in the two tanks and $\mathbf{w}\sim\mathcal{N}(0,10^{-2}I)$ the additive noise. The system parameters are $\tau=0.1$, $\frac{\alpha_{1}}{A_{1}}=1$, $\frac{q_{1}}{A_{1}}=4.5$, $\frac{\alpha_{1}}{A_{2}}=1$, $\frac{\alpha_{2}}{A_{2}}=1$, and $\frac{q_{o}}{A_{2}}=-3$. The domain and initial sets are $X=[1,10]^{2}$ and $X_{0}=[4.75,5.25]^{2}$, respectively.
@@ B.2. Polynomial Systems
where $\mathbf{x}_{k}$ is the room temperature, the ambient temperature $T_{e}=-15^{\circ}$C, and the heater temperature $T_{h}=45^{\circ}$C. Conduction factors are given as $\theta=0.145,\beta=0.06$, and $\nu=-0.0120155\mathbf{x}_{k}+0.8$. The system evolves over the state space $X=[1,50]$ with the initial set $X_{0}=[19.5,20]$, and unsafe regions are $X_{u}=[1,17]\cup[23,50]$. The noise parameter is $\mathbf{w}\sim\mathcal{N}(0,\cdot 10^{-2}I)$.
@@ Appendix D Table Results
Linear 2D Contraction Map 1 2 16.48 1.00 0.00 0.00 0.69 $1.7e^{-2}$ $1.8e^{-2}$ 0.80 0.01 $1.7e^{-2}$ $1.8e^{-2}$ 0.80
4 59.95 1.00 0.00 0.00 1.52 $3.7e^{-3}$ $5.7e^{-3}$ 0.94 0.02 $3.7e^{-3}$ $5.7e^{-3}$ 0.94
8 487.98 1.00 0.00 0.00 12.98 $1.1e^{-3}$ $1.8e^{-3}$ 0.98 0.05 $1.1e^{-3}$ $1.8e^{-3}$ 0.98
12 2157.99 1.00 0.00 0.00 76.57 $1.0e^{-3}$ $1.3e^{-3}$ 0.99 0.23 $1.1e^{-3}$ $1.3e^{-3}$ 0.99
24 OM $\times$ $\times$ $\times$ 2998.25 $8.3e^{-6}$ $9.5e^{-8}$ 0.99 18.74 $9.9e^{-4}$ $1.2e^{-3}$ 0.99
30 OM $\times$ $\times$ $\times$ 9887.87 $4.6e^{-7}$ $7.8e^{-9}$ 0.99 107.47 $9.6e^{-4}$ $1.2e^{-3}$ 0.99
2D Two Tank 10 OM $\times$ $\times$ 289.69 FAILED FAILED 0.00 0.78 $5e^{-7} 2.2$$2e^{-8}$ 1.00
12 OM $\times$ $\times$ $\times$ 1789.52 FAILED FAILED 0.00 1.69 $2.7e^{-9}$ $7.4e^{-10}$ 1.00
14 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 4.49 $2.3e^{-7}$ $6.8e^{-8}$ 1.00
3D Room Temp. 6 OM $\times$ $\times$ $\times$ 1008.67 FAILED FAILED 0.00 10.94 $3.1e^{-7}$ $2.9e^{-8}$ 1.00
8 OM $\times$ $\times$ $\times$ 12897.12 FAILED FAILED 0.00 15.81 $3.1e^{-6}$ $7.5e^{-7}$ 1.00
14 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 9198.78 $1.3e^{-7}$ $3.1e^{-8}$ 1.00
6D Quadrotor 2 OM $\times$ $\times$ $\times$ 9587.45 FAILED FAILED 0.00 0.18 $1.0e^{-1}$ $5.9e^{-4}$ 0.89
4 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 2.57 $9.0e^{-4}$ $1.0e^{-5}$ 1.00
6 OM $\times$ $\times$ $\times$ OM $\times$ $\times$ $\times$ 163.51 $8.8e^{-6}$ $1.7e^{-7}$ 1.00
Polynomial 1D Thermostat 6 0.49 $6.9e^{-9}$ $1.0e^{-9}$ 1.00 0.10 $8.6e^{-9}$ $1.6e^{-5}$ 1.00
8 0.89 $6.9e^{-9}$ $1.9e^{-9}$ 1.00 0.11 $1.5e^{-8}$ $2.8e^{-7}$ 1.00
10 1.13 FAILED FAILED 0.00 0.22 $1.4e^{-8}$ $2.8e^{-7}$ 1.00
2D Oscillator 8 168.79 FAILED FAILED 0.00 3.84 1.00 $2.9e^{-8}$ 0.00
12 1937.42 $5.5e^{-5}$ $8.8e^{-6}$ 1.00 78.89 $1.1e^{-7}$ $1.2e^{-8}$ 1.00
14 4872.54 FAILED FAILED 0.00 144.37 $1.2e^{-7}$ $1.2e^{-8}$ 1.00